OpenStudy (anonymous):

Four more than twice a number is seven more than the number,what is the number 5/3 1 3 9/5

OpenStudy (anonymous):

9/5

OpenStudy (math&ing001):

Let x be the number: 4 + 2x = 7 + x Solve for x to find the number.

OpenStudy (anonymous):

9/5=1.8 multiply it by 2 gives 3.6 add four =3.6+4=7.6

OpenStudy (math&ing001):

@zaibali.qasmi we value the learning process more than the answer it self.